Define the type of given partial differential equation and find the solution.
(y^2+z^2-x^2 )p+2xyq+2zx=0
1
Expert's answer
2020-10-17T13:07:06-0400

Lagrange linear equation


"(y^2+z^2-x^2 )p+2xyq+2zx=0\\
(y^2 + z^2 - x^2)p + 2xyq =- 2zx"


from general form,

"Pp +Qq =R"


"P = y^2 + z^2 -x^2\\
Q = 2xy\\
R =-2zx"


From the auxiliary equation,

"\dfrac{dx}{P}= \dfrac{dy}{Q}= \dfrac{dz}{R}"


"\dfrac{dx}{y^2 +z^2 -x^2}= \dfrac{dy}{2xy}= \dfrac{dz}{-2zx} \quad------(i)"


Taking the last 2 ratios

"\dfrac{dy}{2xy}= \dfrac{dz}{-2zx}"


"\dfrac{dy}{y}= -\dfrac{dz}{z}"


Integrating;

"\log y = -(\log z) + \log c_1\\
\log y = -\log z + \log c_1\\
\log y +\log z = \log c_1\\
\log yz =\log c_1"

"yz = c_1"

"c_1 = yz \quad------(ii)"



Taking the first 2 ratios

"\dfrac{dx}{y^2 +z^2 -x^2}= \dfrac{dy}{2xy}"

but,

"c_1 = yz\\
\therefore z= \dfrac{c_1}{y}"

substituting the value of z into the first 2 ratios,

we have,

"\dfrac{dx}{y^2 +(\frac{c_1}{y})^2 -x^2}= \dfrac{dy}{2xy}"


"2xy\, dx = (y^2 +(\dfrac{c_1}{y})^2 -x^2)dy"


"2xy\, dx - (y^2 +(\dfrac{c_1}{y})^2 -x^2)dy = 0"


"2xy\, dx - (y^2 +\dfrac{c_1^2}{y^2}-x^2)dy = 0 \quad------(iii)"


Integrating equation (iii)


"\large \int2xy\, dx - \int(y^2 +\dfrac{c_1^2}{y^2}-x^2)dy = 0"


"\large x^2y-\dfrac{y^3}{3}+\dfrac{c_1^2}{y} = c_2"


but, "c_1 = yz"

"\therefore\large x^2y-\dfrac{y^3}{3}+yz^2 = c_2 \quad------(iv)"


"\therefore" the general solution is;

"\large f(yz,\,\, x^2y- \dfrac{y^3}{3} +yz^2) = 0"